nolegirl24:

Choose the correct simplification of the expression (3x − 6)(2x^2 − 4x − 5). 6x^3 − 24x^2 + 9x − 30 6x^3 + 9x + 30 6x^3 − 24x^2 + 9x + 30 6x^3 − 24x^2 + 39x + 30

nolegirl24:

Completely confused

dude:

\((3x − 6)(2x^2 − 4x − 5)\) So here you want to distribute

dude:



nolegirl24:

Okay okay hold on imma write it down

dude:

Okay

nolegirl24:

Okay done

dude:

\[(\color{red}{3x}\color{blue}{-6})(2x^2-4x-5)=\] \[(\color{red}{3x}\times 2x^2)+(\color{red}{3x}\times-4x)+(\color{red}{3x}\times-5)+(\color{blue}{-6}\times2x^2)+(\color{blue}{-6}\times-4x)+(\color{blue}{-6}\times-5) \]

dude:

Latex cut off RIP

dude:

\[(\color{red}{3x}\color{blue}{-6})(2x^2-4x-5)=\] \[(\color{red}{3x}\times 2x^2)+(\color{red}{3x}\times-4x)+(\color{red}{3x}\times-5)+\\(\color{blue}{-6}\times2x^2)+(\color{blue}{-6}\times-4x)+(\color{blue}{-6}\times-5) \]

nolegirl24:

Got it just writing it down :)

nolegirl24:

Okie done

dude:

Well just simplify those

dude:

Do you know how to do that?

nolegirl24:

Multiply the ones in () and then add?

nolegirl24:

Oh wait idk shoot o.o

dude:

Yes multiply the ones in parenthesis

nolegirl24:

What about the ones that are like 3x * -5

dude:

Multiply the number only, so it would be -15x

nolegirl24:

So is this right? \[6x^{2} + 12x + 15 - 12x ^{2} + 24x + 30\]

dude:

When multiplying exponents you add the exponent ex. \(x\times x^2=x^3\)

nolegirl24:

so 6x^2 would become 6x^3?

dude:

Right

nolegirl24:

Alright lemme fix

nolegirl24:

Okay done

dude:

Okay, what do you have?

nolegirl24:

\[6x^3 - 12x + 15x - 12x^3 + 24x + 30\]

dude:

Almost \((\color{red}{3x}\times 2x^2)=6x^3 ~~\\ +(\color{red}{3x}\times-4x)=-12x^2\\ +(\color{red}{3x}\times-5)=-15x\\ +(\color{blue}{-6}\times2x^2)=-12x^3\\+ (\color{blue}{-6}\times-4x)=24x\\ +(\color{blue}{-6}\times-5)=30\)

nolegirl24:

ohh

dude:

The power adding applies to all exponents \(x\times x=x^2\) \(x\times x^3=x^4\) \(number~\times x=numberx\)

nolegirl24:

time to rewrite

nolegirl24:

Alright i think i got it now

dude:

Okay

nolegirl24:

Idk what to do now ;-;

dude:

Add the numbers that have the same power of x ex. \(2x+5+3x+9\) = \(5x+14\) This applies to app powers!

nolegirl24:

So like 6x³ - 12x^3 ?

nolegirl24:

would the exponets add?

dude:

Yes, add the co-efficients

dude:

No exponents stay the same

nolegirl24:

So like this? 6x^3 + 12x^3 = 18x^3

dude:

Right

nolegirl24:

Okay so this is what i got. \[18x^3 - 12x^2 - 39x + 30\]

dude:

Ah! i goofed previously! \(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Almost \((\color{red}{3x}\times 2x^2)=6x^3 ~~\\ +(\color{red}{3x}\times-4x)=-12x^2\\ +(\color{red}{3x}\times-5)=-15x\\ +(\color{blue}{-6}\times2x^2)=-12x^3\\+ (\color{blue}{-6}\times-4x)=24x\\ +(\color{blue}{-6}\times-5)=30\) \(\color{#0cbb34}{\text{End of Quote}}\) \(+(\color{blue}{-6}\times2x^2)=-12x^3\) Correction: \(+(\color{blue}{-6}\times2x^2)=-12x^2\)

nolegirl24:

Now i am confused...............

dude:

I made a mistake, in \(+(\color{blue}{-6}\times2x^2)=-12x^3\) I told you about the exponents rule right? I accidentally added an exponent here -6 doesn't have an x, so the \(x^2\) stays the same

dude:

This was just a mistake from my end

nolegirl24:

so it would be 18x^2?

dude:

\(6x^3\color{red}{-12x^2}-15x-12x^3+24x+30\) is your equation only the red segment is the change

nolegirl24:

Alright i fixed it

nolegirl24:

So now i do the same that i just did with adding them?

dude:

Yes

nolegirl24:

so 6x^3 + -12x^3?

dude:

\(-12x^3\) is \(-12x^2\) when I corrected myself So \(6x^3\) stays like that \(6x^3-12x^2-15x-12x^2+24x+30\)

nolegirl24:

Oh yeah dang it my bad

nolegirl24:

\[6x^3 - 12x^2 - 12x^2 - 15x + 24x + 30\] so like this^^^^

dude:

Right

nolegirl24:

both of the 12x^2 would cancel out?

dude:

They are both negative, they would add up

nolegirl24:

so it would end up being 24x^2?

dude:

-24x^2, yes, keep the sign

nolegirl24:

Okay

nolegirl24:

\[6x^3 - 24x^2 - 9x +30\] Like this?

dude:

Close, when adding −15x+24x You want to keep the sign of the greater number

nolegirl24:

so positive

dude:

Right

nolegirl24:

I got it now

nolegirl24:

So the answer would be C?

dude:

Yep

nolegirl24:

THANK YOUUUUUUUU

nolegirl24:

Iḿma end up writing it all back out and neater so i can get it again :D. Thanks again!

dude:

Okay, sure!